# Theoretical Guarantees for the Subspace-Constrained Tyler’s Estimator

### v1

@@ 1 Introduction
The subspace-constrained Tyler’s estimator (STE) is a recent method for robust subspace recovery (RSR), which was recently suggested in [2]. In particular, [2] explains state two theorems for STE. While [2] addresses the computer vision community, we believe that the proofs of these theorems are of independent interest to some mathematicians and thus include them in this preprint, which will eventually evolve into a complete manuscript.
@@ 1.1 Notation
We use bold upper and lower case letters for matrices and column vectors, respectively. Let ${\mathbf{I}}_{k}$ denote the identity matrix in $\mathbb{R}^{k\times k}$, where if $k$ is obvious we just write ${\mathbf{I}}$. For a matrix ${\mathbf{A}}$, we denote by $\operatorname{tr}({\mathbf{A}})$ and $\operatorname{Im}({\mathbf{A}})$ the trace and image (i.e., column space) of ${\mathbf{A}}$. We denote by $S_{+}(D)$ and $S_{++}(D)$ the sets of positive semidefinite and positive definite matrices in $\mathbb{R}^{D\times D}$, respectively. We denote by $O(D,d)$ the set of semi-orthogonal $\mathbb{R}^{D\times d}$, i.e., ${\mathbf{U}}\in O(D,d)$ if and only if ${\mathbf{U}}\in\mathbb{R}^{D\times d}$ and ${\mathbf{U}}^{T}{\mathbf{U}}={\mathbf{I}}_{d}$. For a $d$-subspace $L$, we denote by ${\mathbf{P}}_{L}$ the $D\times D$ matrix representing the orthogonal projector onto $L$. This matrix is symmetric and satisfies ${\mathbf{P}}^{2}_{L}={\mathbf{P}}_{L}$ and $\operatorname{Im}({\mathbf{P}}_{L})=L$. We also arbitrarily fix ${\mathbf{U}}_{L}$ in $O(D,d)$ such that ${\mathbf{U}}_{L}{\mathbf{U}}_{L}^{T}={\mathbf{P}}_{L}$ (such ${\mathbf{U}}_{L}$ is determined up to right multiplication by an orthogonal matrix in $O(d,d)$). Throughout the paper, ${\mathcal{X}}=\{{\bm{x}}_{i}\}_{i=1}^{N}\subset{\mathbb{R}}^{D}$ is assumed to be a given centered dataset, that is, $\sum_{i=1}^{N}{\bm{x}}_{i}={\mathbf{0}}$. We refer to linear $d$-dimensional subspaces as $d$-subspaces.
@@ 1.2 Review of the STE Algorithm
Algorithm 1 STE: Subspace-Constrained Tyler’s Estimator [2]
@@
8:     $\widetilde{{\mathbf{S}}}^{(k)}\leftarrow\operatorname{diag}(\sigma_{1},\ldots,\sigma_{d},\gamma\sigma_{d+1,D},\ldots,\gamma\sigma_{d+1,D})$,
@@
12:  $L$ = first $d$ column of ${\mathbf{U}}^{(k)}$
@@ 1.3 Review of the Stated Theory
$\kappa_{2}=\frac{\sigma_{1}\Big{(}{\mathbf{\Sigma}}^{(0)}_{L_{*}^{\perp},L_{*}^{\perp}}\Big{)}}{\sigma_{D}({\mathbf{\Sigma}}^{(0)})}\ \text{ and }\ \kappa_{in}=\frac{\sigma_{1}({\mathbf{\Sigma}}_{in,*})}{\sigma_{d}({\mathbf{\Sigma}}_{in,*})},$
where we note that $\kappa_{in}$ is analogous to the one in (25) of [8], where we replace the sample covariance by the TME estimator. An analog to the alignment of outliers statistic [7, 9] for STE is
@@
The exact technical requirement on $C$ is specified in our proof. In general, the larger the RHS of (3), the more restricted the choice of ${\mathbf{\Sigma}}^{(0)}$ is. In particular, when $\kappa_{1}=\infty$, the definition of $\kappa_{1}$ implies that $\operatorname{Im}({\mathbf{\Sigma}}^{(0)})=L_{*}$, so the subspace is already recovered by the initial estimate. Therefore, reducing the lower bound of $\kappa_{1}$ may allow some flexibility, so a marginally suboptimal initialization could still work out. In §2 we show that under the asymptotic generalized haystak model, Assumption 3 can be interpreted as an upper bound on the largest principal angle between the initial and ground truth subspaces.
Generic Theory: The next theorem suggests that under assumptions 1 and 2, STE nicely converges to an estimator that recovers $L_{*}$. The main significance of this theory is that its assumptions can allow DS-SNR lower than 1 for special instances of datasets (for which the assumptions hold), unlike the general recovery theories of [13] and [3].
@@
Under assumptions 1 and 2, the sequence ${\mathbf{\Sigma}}^{(k)}$ generated by STE converges to ${\mathbf{U}}_{L_{*}}{\mathbf{\Sigma}}_{in,*}{\mathbf{U}}_{L_{*}}^{\top}$, the TME solution for the set of inliers ${\mathcal{X}}_{in}$. In addition, let $L^{(k)}$ be the subspace spanned by the top $d$ eigenvectors of ${\mathbf{\Sigma}}^{(k)}$, then the angle between $L^{(k)}$ and $L_{*}$, $\angle(L^{(k)},L_{*})=\cos^{-1}(\|{\mathbf{U}}_{L^{(k)}}^{T}{\mathbf{U}}_{L_{*}}\|)$, converges $r$-linearly to zero.
@@
Theory under a probabilistic model: We show that under a common probabilistic model, the assumptions of Theorem 1, where ${\mathbf{\Sigma}}^{(0)}$ is obtained by TME, hold. Moreover, we show that STE (initialized by TME) can recover the correct subspace in situations with DS-SNR$<1$, whereas TME cannot recover the underlying subspace in such cases. We follow [9] and study the Generalized Haystack Model, though for simplicity, we assume Gaussian instead of sub-Gaussian distributions. We assume $n_{1}$ inliers i.i.d. sampled from a Gaussian distribution $N(0,{\mathbf{\Sigma}}^{(in)}/d)$, where ${\mathbf{\Sigma}}^{(in)}\in S_{+}(D)$ and $L_{*}=\operatorname{Im}({\mathbf{\Sigma}}^{(in)})$ (so ${\mathbf{\Sigma}}^{(in)}$ has $d$ nonzero eigenvalues), and $n_{0}$ outliers are i.i.d. sampled from a Gaussian distribution $N(0,{\mathbf{\Sigma}}^{(out)}/D)$, where ${\mathbf{\Sigma}}^{(out)}/D\in S_{++}(D)$. We define the following condition numbers of inliers (in $L_{*}$) and outliers:
@@
Clearly, Assumption 1 holds under this model. Our next theorem shows that Assumption 3 holds under this model when the initial estimate ${\mathbf{\Sigma}}^{(out)}$ for STE is obtained by TME. It also shows that in this case STE can solve the RSR problem even when DS-SNR$<1$, unlike TME. For simplicity, we formulate the theory for the asymptotic case, where $N\rightarrow\infty$ and the theorem holds almost surely. It is possible to formulate it for a very large $N$ with high probability, but it requires stating complicated constants depending on various parameters.
@@ 3.1.2 Preliminaries: Some Useful Results
The proof of part (a) follows from the Courant-Fischer min-max theorem for the eigenvalues [10, Theorem 1.3.2]. For completeness we verify this as follows:
@@ 3.1.4 Proof of (30)
The second equality is the second extremal partial trace formula in [10, Proposition 1.3.4] and the second inequality applies the trace bound in (18). The third equality is because the inliers lie on $L_{*}$ and thus ${\mathbf{U}}_{L_{*}^{\perp}}^{\top}{\mathbf{\Sigma}}_{+,in}{\mathbf{U}}_{L_{*}^{\perp}}={\mathbf{0}}$.
@@ 3.2.1 Proof of Lemma 10
By the geodesic convexity of $F$ [13, 12], the geometric mean of ${{\mathbf{\Sigma}}}^{(TME)}$ and ${\mathbf{\Sigma}}^{(TME)\,\star}$ should also be the minimizer of $F$. However, for any ${\mathbf{\Sigma}}$, the geodesic mean of ${\mathbf{\Sigma}}$ and ${\mathbf{\Sigma}}^{\star}$, denoted by ${\bf{M}}$, should be “block diagonal” in the sense that ${\bf{M}}_{L_{*},L_{*}^{\perp}}={\mathbf{0}}$ and ${\bf{M}}_{L_{*}^{\perp},L_{*}}={\mathbf{0}}$. One can obtain this property by applying the geometric mean algorithm in [4, (4)], which is block diagonal in every iteration. Since the TME solution is unique (up to scaling), the TME solution is block diagonal.

### v2

@@ 1 Introduction
The subspace-constrained Tyler’s estimator (STE) is a recent method for robust subspace recovery (RSR), which was recently suggested in [12]. In particular, [12] states two theorems for STE. While [12] addresses the computer vision community, we believe that the proofs of these theorems are of independent interest to some mathematicians and thus include them in this preprint, which will eventually evolve into a complete manuscript.
@@ 1.1 Notation
We use bold upper and lower case letters for matrices and column vectors, respectively. Let ${\mathbf{I}}_{k}$ denote the identity matrix in $\mathbb{R}^{k\times k}$, where if $k$ is obvious we just write ${\mathbf{I}}$. For a matrix ${\mathbf{A}}$, we denote by $\operatorname{tr}({\mathbf{A}})$ and $\operatorname{Im}({\mathbf{A}})$ the trace and image (i.e., column space) of ${\mathbf{A}}$. We denote by $S_{+}(D)$ and $S_{++}(D)$ the sets of positive semidefinite and definite matrices in $\mathbb{R}^{D\times D}$, respectively. We denote by $O(D,d)$ the set of semi-orthogonal ${D\times d}$ matrices, i.e., ${\mathbf{U}}\in O(D,d)$ if and only if ${\mathbf{U}}\in\mathbb{R}^{D\times d}$ and ${\mathbf{U}}^{\top}{\mathbf{U}}={\mathbf{I}}_{d}$. We refer to linear $d$-dimensional subspaces as $d$-subspaces. For a $d$-subspace $L$, we denote by ${\mathbf{P}}_{L}$ the $D\times D$ matrix representing the orthogonal projector onto $L$. We also arbitrarily fix ${\mathbf{U}}_{L}$ in $O(D,d)$ such that ${\mathbf{U}}_{L}{\mathbf{U}}_{L}^{\top}={\mathbf{P}}_{L}$ (such ${\mathbf{U}}_{L}$ is determined up to right multiplication by an orthogonal matrix in $O(d,d)$). Throughout the paper, ${\mathcal{X}}=\{{\bm{x}}_{i}\}_{i=1}^{N}\subset{\mathbb{R}}^{D}$ is assumed to be a given centered dataset, that is, $\sum_{i=1}^{N}{\bm{x}}_{i}={\mathbf{0}}$.
@@ 1.2 Review of the STE Algorithm
Algorithm 1 STE: Subspace-Constrained Tyler’s Estimator
@@
8:     $\widetilde{{\mathbf{S}}}^{(k)}\leftarrow\operatorname{diag}(\sigma_{1},\ldots,\sigma_{d},\gamma\cdot\sigma_{d+1,D},\ldots,\gamma\cdot\sigma_{d+1,D})$,
@@
12:  $L$ = Span of the first $d$ columns of ${\mathbf{U}}^{(k)}$
@@ 1.3 Review of the Stated Theory
$\kappa_{2}=\frac{\sigma_{1}\Big{(}{\mathbf{\Sigma}}^{(0)}_{L_{*}^{\perp},L_{*}^{\perp}}\Big{)}}{\sigma_{D}({\mathbf{\Sigma}}^{(0)})}\ \text{ and }\ \kappa_{in}=\frac{\sigma_{1}({\mathbf{\Sigma}}_{in,*})}{\sigma_{d}({\mathbf{\Sigma}}_{in,*})}.$
We note that $\kappa_{in}$ is analogous to the one in (25) of [7], where we replace the sample covariance by the TME estimator. An analog to the alignment of outliers statistic [6, 8] for STE is
@@
The exact technical requirement on $C$ is specified in our proof. In general, the larger the RHS of (3), the more restricted the choice of ${\mathbf{\Sigma}}^{(0)}$ is. In particular, when $\kappa_{1}=\infty$, the definition of $\kappa_{1}$ implies that $\operatorname{Im}({\mathbf{\Sigma}}^{(0)})=L_{*}$, so the subspace is already recovered by the initial estimate. Therefore, reducing the lower bound of $\kappa_{1}$ may allow some flexibility, so a marginally suboptimal initialization could still work out. In §2 we show that under the asymptotic generalized haystack model, Assumption 3 can be interpreted as an upper bound on the largest principal angle between the initial and ground truth subspaces.
Generic Theory: The next theorem suggests that under assumptions 1-3, STE nicely converges to an estimator that recovers $L_{*}$. The main significance of this theory is that its assumptions can allow DS-SNR lower than 1 for special instances of datasets (for which the assumptions hold), unlike the general recovery theories of [2] and [13]. The main significance of this theory is that its assumptions can allow DS-SNR lower than 1 for special instances of datasets (for which the assumptions hold), unlike the general recovery theories of [13] and [2].
@@
Under assumptions 1-3, the sequence ${\mathbf{\Sigma}}^{(k)}$ generated by STE converges to ${\mathbf{U}}_{L_{*}}{\mathbf{\Sigma}}_{in,*}{\mathbf{U}}_{L_{*}}^{\top}$, the TME solution for the set of inliers ${\mathcal{X}}_{in}$. In addition, let $L^{(k)}$ be the subspace spanned by the top $d$ eigenvectors of ${\mathbf{\Sigma}}^{(k)}$, then the angle between $L^{(k)}$ and $L_{*}$, $\angle(L^{(k)},L_{*})=\cos^{-1}(\|{\mathbf{U}}_{L^{(k)}}^{\top}{\mathbf{U}}_{L_{*}}\|)$, converges $r$-linearly to zero.
@@
Theory under a probabilistic model: We show that under a common probabilistic model, the assumptions of Theorem 1, where ${\mathbf{\Sigma}}^{(0)}$ is obtained by TME, hold. Moreover, we show that STE (initialized by TME) can recover the correct subspace in situations with DS-SNR$<1$, whereas TME cannot recover the underlying subspace in such cases. We follow [8] and study the Generalized Haystack Model, though for simplicity, we assume Gaussian instead of sub-Gaussian distributions and an asymptotic setting. We assume $n_{1}$ inliers i.i.d. sampled from a Gaussian distribution $N(0,{\mathbf{\Sigma}}^{(in)}/d)$, where ${\mathbf{\Sigma}}^{(in)}\in S_{+}(D)$ and $L_{*}=\operatorname{Im}({\mathbf{\Sigma}}^{(in)})$ (so ${\mathbf{\Sigma}}^{(in)}$ has $d$ nonzero eigenvalues), and $n_{0}$ outliers are i.i.d. sampled from a Gaussian distribution $N(0,{\mathbf{\Sigma}}^{(out)}/D)$, where ${\mathbf{\Sigma}}^{(out)}/D\in S_{++}(D)$. We define the following condition numbers of inliers (in $L_{*}$) and outliers:
@@
Clearly, Assumption 1 holds under this model, and Assumption 2 constrains some of its parameters. Our next theorem shows that Assumption 3 holds under this model when the initial estimate ${\mathbf{\Sigma}}^{(out)}$ for STE is obtained by TME. It also shows that in this case STE can solve the RSR problem even when DS-SNR$<1$, unlike TME. For simplicity, we formulate the theory for the asymptotic case, where $N\rightarrow\infty$ and the theorem holds almost surely. It is possible to formulate it for a very large $N$ with high probability, but it requires stating complicated constants depending on various parameters.
@@ 3.1.2 Preliminaries: Some Useful Results
The proof of part (a) follows from the Courant-Fischer min-max theorem for the eigenvalues [9, Theorem 1.3.2]. For completeness we verify this as follows:
@@ 3.1.4 Proof of (30)
The second equality is the second extremal partial trace formula in [9, Proposition 1.3.4] and the second inequality applies the trace bound in (18). The third equality is because the inliers lie on $L_{*}$ and thus ${\mathbf{U}}_{L_{*}^{\perp}}^{\top}{\mathbf{\Sigma}}_{+,in}{\mathbf{U}}_{L_{*}^{\perp}}={\mathbf{0}}$.
@@ 3.2.1 Proof of Lemma 10
By the geodesic convexity of $F$ [13, 11], the geometric mean of ${{\mathbf{\Sigma}}}^{(TME)}$ and ${\mathbf{\Sigma}}^{(TME)\,\star}$ should also be the minimizer of $F$. However, for any ${\mathbf{\Sigma}}$, the geodesic mean of ${\mathbf{\Sigma}}$ and ${\mathbf{\Sigma}}^{\star}$, denoted by ${\bf{M}}$, should be “block diagonal” in the sense that ${\bf{M}}_{L_{*},L_{*}^{\perp}}={\mathbf{0}}$ and ${\bf{M}}_{L_{*}^{\perp},L_{*}}={\mathbf{0}}$. One can obtain this property by applying the geometric mean algorithm in [3, (4)], which is block diagonal in every iteration. Since the TME solution is unique (up to scaling), the TME solution is block diagonal.